# Comfortability of quantum walks on embedded graphs on surfaces

### v1

@@ 6.1 The scattering matrix represented by faces
$\boldsymbol{\beta}_{out}=S\boldsymbol{\alpha}_{in},$
which is determined by the rotation system $G(\rho,\tau)$ and $\delta X$, and independent of $\boldsymbol{\alpha}_{in}$, $\boldsymbol{\beta}_{out}$. It is shown in [8] that such a matrix $S$ exists and is unitary. However its explicit expression is up to the individual setting.
@@
for any $j\in\{0,1,\dots,\kappa-1\}$ in the modulus of $\kappa$. Here if $\xi_{j}\in\delta A_{is}$, $\xi_{j}$ represents the quay arcs $(\xi_{j}^{+},\xi_{j}^{-})$. Let $\delta F\subset F$ be the set of (extended) facial walk passing through a boundary vertex $\delta X$ in $\tilde{G}$. For any facial walk $f$, there exists a chiral facial walk $f^{*}$ defined by going around the opposite direction on the opposite sheet. We introduce the following useful lemma to consider the scattering.
@@
$\psi_{\infty}(\xi^{\prime}_{j+1})=(-1)^{\tau(e_{j})}\omega\;\psi_{\infty}(\xi_{j}^{\prime\prime}),\;(j=0,1,\dots,\kappa-1)$
where if $\xi_{j+1}\in\delta A_{is}$, then $\xi_{j+1}^{\prime}=\xi_{j+1}^{+}$, otherwise $\xi_{j+1}$; if $\xi_{j}\in\delta A_{is}$, then $\xi_{j}^{\prime\prime}=\xi_{j}^{-}$, otherwise $\xi_{j}^{\prime\prime}=\xi_{j}$.
@@
For any bridge arcs $e\in A_{br}$, (of course) the vertices $o(e)$ and $t(e)$ are connected by the bridge arcs $e^{R}_{1}:=e$ and $e^{L}_{1}:=\bar{e}$, moreover the island arcs $e_{0}^{R}:=\mathrm{is}(e)$ and $e_{0}^{L}:=\mathrm{is}^{\sharp}(e)$ are connected to $o(e)$, while the island arcs $e_{2}^{R}:=\mathrm{is}^{\sharp}(\bar{e})$ and $e_{2}^{L}:=\mathrm{is}(\bar{e})$ are connected to $o(t)$. Then the blow up graph has locally a path structure with $2$ vertices and “$6$” symmetric arcs. Let us see this fact leads a transfer matrix discussed in the spectral analysis on the discrete-time quantum walks on the one-dimensional lattice and gives the conclusion. Let us set $\psi:=\psi_{\infty}$. The point to get the transfer matrix is to align the “subscriptions” for each vector: the local eigenequation is rewritten by
@@
where $\omega^{\prime}=(-1)^{\tau(e)}\omega$. Here the second equality is obtained by the properties of the unitarity matrix $C$, for examples, $a=-\omega^{\prime}\bar{d^{\prime}}$, $b=\omega^{\prime}\bar{c^{\prime}}$, $|d^{\prime}|^{2}+|c^{\prime}|^{2}=1$. Then $d\in\mathbb{R}$ if and only if $\psi(e_{2}^{R})=\omega^{\prime}\psi(e_{0}^{R})$ and $\psi(e_{0}^{L})=\omega^{\prime}\psi(e_{2}^{L})$, which is the desired conclusion. ∎
For a facial walk $f=(\xi_{0},e_{0},\dots,\xi_{\kappa-1},e_{\kappa-1})\in\delta F$, set $f\cap\delta X=\{x_{k_{0}},\dots,x_{k_{q-1}}\}$ and $f\cap\delta A_{is}=\{\xi_{k_{0}},\dots,\xi_{k_{q}}\}$ with $k_{0}<k_{1}<\cdots<k_{q-1}$ in the modulus of $q$. Here $1\leq q\leq\kappa$. Note that the $q$ tails interfere with the facial walk $f$. The tail originating the boundary $x_{k_{i}}$ is denoted by $\mathrm{Tail}_{i}$. Put
$\bowtie_{f}(\ell,m):=\int_{x_{k_{\ell}}}^{x_{k_{m}}}\tau\text{ and }\mathrm{dist}_{f}(\ell,m):=k_{\ell}-k_{m},$
which are the parity of number of type-$1$ edge between $x_{m}$ and $x_{\ell}$ , and the distance between the boundary vertices $k_{m}$ and $k_{\ell}$ along the facial walk $f$. To simplify the notation, let us put $f\cap\delta X=\{0,1,\dots,q-1\}$. We set the identity matrix, the weighted cyclic permutation matrix on $\mathbb{C}^{\{0,\dots,q-1\}}$ by
@@
From this expression, if $f\cap\delta X=\{k_{0},\dots,k_{q-1}\}$, we have
$(S_{f})_{k_{j},k_{i}}=\begin{cases}bc\frac{a^{j-i-1}\omega^{{\rm dist}_{f}(k_{j},k_{i})}}{1-a^{|f\cap\delta X|}}\times(-1)^{\bowtie_{f}(k_{j},k_{i})}&\text{: $i\neq j$,}\\ \\ bc\frac{a^{|f\cap\delta X|-1}\omega^{|f|}}{1-a^{|f\cap\delta X|}}+d&\text{: $i=j$.}\end{cases}$
@@
Pick up a facial walk $f=(\xi_{0},e_{0},\dots,\xi_{\kappa-1},e_{\kappa-1})\in\delta F$, and set $f\cap\delta X=\{x_{j_{0}},\dots,x_{j_{q-1}}\}$ and $f\cap\delta A_{is}=\{\xi_{k_{0}},\dots,\xi_{k_{q}}\}$. Here $1\leq q\leq\kappa$. Note that the $q$ tails interfere with the facial walk $f$. The tail originating the boundary $x_{k_{i}}$ is denoted by $\mathrm{Tail}_{i}$. Let us give the inflow $1$ from the $\mathrm{Tail}_{i}$, and consider the outflow to the tail $\mathrm{Tail}_{j}$ ($i,j\in\{0,1,\dots,\kappa-1\}$). Let us see that the $(j,i)$ element of the scattering matrix $S$ can be calculated according to the number of “nights” that a quantum walk stays at the face $f$ with the boundary vertices $x_{k_{i}}$ and $x_{k_{j}}$ from the time the quantum walk enters at entrance $i$ to the time the quantum walk leaves at exit $j$.
Consider the case for $i=0$. Let $t^{in}_{j}\in\delta A_{pr}^{+}$ and $t^{out}_{j}\in\delta A_{pr}^{-}$ be the arcs of $\mathrm{Tail}_{j}$ originating the vertex $x_{k_{j}}\in\delta X$. Let “the day trip walk” from $\mathrm{Tail}_{0}$ to $\mathrm{Tail}_{j}$ be defined by
@@
$(t_{0}^{in},\boldsymbol{w}_{0},\overbrace{\xi_{k_{j}}^{+},e_{k_{j}},\dots,\xi_{k_{j+1}}^{-},\xi_{k_{j+1}}^{+},e_{k_{j+1}},\dots,\xi_{k_{j-1}}^{-},\xi_{k_{j-1}}^{+},e_{k_{j-1}},\dots,\xi_{k_{j}}^{-},}^{\text{$r$ times (nights)}}t_{j}^{out})$
@@
the weights associated with moving from the tail $t_{0}^{in}\in\delta A_{pr}^{+}$ to the quay $\xi_{k_{0}}^{+}\in A_{is}$, and from the quay $\xi_{k_{j}}^{-}\in A_{is}$ to the tail $t_{j}^{out}\in\delta A_{pr}^{-}$ are $c$ and $b$, respectively; the weight associated with moving from $\xi_{k_{s}}^{-}\in A_{is}$ to $\xi_{k_{s}}^{+}\in A_{is}$ is $a$. Remark that the weight of the closed path starting from $\xi_{{k_{0}}}^{+}$ and returning back to $\xi_{k_{0}}^{+}$ along the boundary face $f$ is $a^{q}\omega^{\kappa}$ since $(-1)^{\int_{f}\tau}=1$. Then set $\boldsymbol{s}^{(r)}(\cdot,0)\in\mathbb{C}^{\{0,1,\dots,q-1\}}$ ($r=0,1,\dots$) as the weight of $r$-night walk in the stationary state by $\boldsymbol{s}^{(r)}(j,0)=(a^{q}\omega^{\kappa})^{r}\times\boldsymbol{s}^{(0)}(j,0)$, $(r=1,\dots,\kappa-1)$ with
$\boldsymbol{s}^{(0)}(j,0)=\begin{cases}b\;c\;a^{j-1}\;\omega^{k_{j}-k_{0}}\;(-1)^{\int_{x_{k_{0}}}^{x_{k_{j}}}\tau}&\text{: $j\neq 0$}\\ b\;c\;a^{q-1}\;\omega^{\kappa}\;&\text{: $j=0$}\end{cases}$
@@
$(S)_{k_{j},k_{0}}$ $=\sum_{r=0}^{\infty}\boldsymbol{s}^{(r)}(j,0)=\begin{cases}bc\frac{a^{j-1}\omega^{k_{j}-k_{0}}}{1-a^{q}\omega^{\kappa}}\times(-1)^{\int_{x_{k_{0}}}^{x_{k_{j}}}\tau}&\text{: $j\neq 0$}\\ \\ bc\frac{a^{q-1}\omega^{\kappa}}{1-a^{q}\omega^{\kappa}}+d&\text{: $j=0$}\end{cases}$
@@
$(S)_{k_{j},k_{i}}=\begin{cases}bc\frac{a^{j-i-1}\omega^{k_{j}-k_{i}}}{1-a^{q}\omega^{\kappa}}\times(-1)^{\int_{x_{k_{i}}}^{x_{k_{j}}}\tau}&\text{: $i\neq j$}\\ \\ bc\frac{a^{q-1}\omega^{\kappa}}{1-a^{q}\omega^{\kappa}}+d&\text{: $i=j$,}\end{cases}$ (6.25)
@@
Let $(G,\rho_{1},\tau_{1})$ and $(G,\rho_{2},\tau_{2})$ be the rotation systems which are isomorphic to each other. Let $S_{1}$ and $S_{2}$ be the scattering matrices of those resulting embbedings, respectively.
@@
Let $S_{1}$ and $S_{2}$ be the above. Then we have
@@ 6.2 Detection of the orientability
Let us estimate whether the underlying closed surface is orientable or not by observing the outflow of the internal graph to an inflow. The following theorem may be useful for such an estimation.
@@
Under Assumption 1 with $a>0$, the underlying surface is orientable if and only if for any $x,y\in X^{\tau}$ with $x\neq y$, the signature of the element of scattering matrix
$\mathrm{sgn}((S)_{e_{x},e_{y}})$
is invariant for any $(e_{x},e_{y})\in\{e,e^{\prime}\in A^{\tau};\;|\;(S)_{e,e^{\prime}}\neq 0,\;t(e)=x,t(e^{\prime})=y\}$.
@@
$(S)_{e,e^{\prime}}/|S_{e,e^{\prime}}|=\begin{cases}(-1)^{\int_{t(e)}^{t(e^{\prime})}\tau}&\text{: $e$ and $e^{\prime}$ are included in the same face, }\\ 0&\text{: otherwise.}\end{cases}$
Here $\int_{x}^{y}\tau:=\sum_{e^{\prime\prime}\in p}\tau(e^{\prime\prime})$ with a path $p$ satisfying $o(p)=x$ and $t(p)=y$. Note that Lemma 4.1 implies that $\int_{x}^{y}\tau$ is independent of the choice of path if and only if the underlying surface is orientable. ∎
This means that once a vertex is found where there exist different signatures among the outflows from it, then the underlying closed surface must be unorientable.

### v2

@@ 6.1 The scattering matrix represented by faces
$\boldsymbol{\beta}_{out}=S\boldsymbol{\alpha}_{in},$ (6.25)
which is determined by the rotation system $G(\rho,\tau)$ and $\delta X$, and independent of $\boldsymbol{\alpha}_{in}$, $\boldsymbol{\beta}_{out}$. It is shown in [10] that such a matrix $S$ exists and is unitary, but its explicit expression is up to the individual setting. Here we characterize the scattering matrix $S$ by the geometric information.
@@
for any $j\in\{0,1,\dots,\kappa-1\}$ in the modulus of $\kappa$. We remark that $\xi_{j}$ represents the sequence of two quay arcs $\xi_{j}^{-}$ and $\xi_{j}^{+}$ if $\xi_{j}\in\delta A_{is}$. Let $\delta F\subset F$ be the set of (extended) facial walk passing through a boundary vertex $\delta X$ in $\tilde{G}$. For any facial walk $f$, there exists a chiral facial walk $f^{*}$ defined by going around the opposite direction on the opposite sheet. We introduce the following useful lemma to consider the scattering.
@@
$\psi_{\infty}(\xi^{\prime}_{j+1})=(-1)^{\tau(e_{j})}\omega\;\psi_{\infty}(\xi_{j}^{\prime\prime}),\;(j=0,1,\dots,\kappa-1)$ (6.28)
where if $\xi_{j+1}\in\delta A_{is}$, then $\xi_{j+1}^{\prime}=\xi_{j+1}^{-}$, otherwise $\xi_{j+1}$; if $\xi_{j}\in\delta A_{is}$, then $\xi_{j}^{\prime\prime}=\xi_{j}^{+}$, otherwise $\xi_{j}^{\prime\prime}=\xi_{j}$.
@@
For any bridge arcs $e\in A_{br}$, (of course) the vertices $o(e)$ and $t(e)$ are connected by the bridge arcs $e^{R}_{1}:=e$ and $e^{L}_{1}:=\bar{e}$, moreover the island arcs $e_{0}^{R}:=\mathrm{is}(e)$ and $e_{0}^{L}:=\mathrm{is}^{\sharp}(e)$ are connected to $o(e)$, while the island arcs $e_{2}^{R}:=\mathrm{is}^{\sharp}(\bar{e})$ and $e_{2}^{L}:=\mathrm{is}(\bar{e})$ are connected to $o(t)$. Let us set $\psi:=\psi_{\infty}$. The local scatterings at $o(e)$ and $t(e)$ imply
@@
where $\omega^{\prime}=(-1)^{\tau(e)}\omega$. Here the second equality is obtained by the properties of the unitarity matrix $C$, for examples,
$a=-\omega^{\prime}\bar{d^{\prime}},\;b=\omega^{\prime}\bar{c^{\prime}},$
$|a|^{2}+|b|^{2}=|d^{\prime}|^{2}+|c^{\prime}|^{2}=1$ (6.30)
and so on. Then $d\in\mathbb{R}$ if and only if $\psi(e_{2}^{R})=\omega^{\prime}\psi(e_{0}^{R})$ and $\psi(e_{0}^{L})=\omega^{\prime}\psi(e_{2}^{L})$, which is the desired conclusion. ∎
The local path structures around the bridges in the blow-up graph lead transfer matrices discussed in the spectral analysis on the discrete-time quantum walks on the one-dimensional lattice and gives the conclusion.
For a facial walk $f=(\xi_{0},e_{0},\dots,\xi_{\kappa-1},e_{\kappa-1})\in\delta F$, set $f\cap\delta A_{is}=\{\xi_{k_{0}},\dots,\xi_{k_{q-1}}\}$ and $f\cap\delta X:=\{x_{k_{0}},\dots,x_{k_{q-1}}\}$, where $x_{k_{j}}=t(\xi_{k_{j}}^{-})$ with $k_{0}<k_{1}<\cdots<k_{q-1}$ in the modulus of $q$. Here $1\leq q\leq\kappa$. (See Figure 11. ) Note that the $q$ tails interfere with the facial walk $f$. Put
$\bowtie_{f}(\ell,m):=\int_{x_{k_{\ell}}}^{x_{k_{m}}}\tau\text{ and }\ \mathrm{dist}_{f}(\ell,m):=k_{\ell}-k_{m},$ (6.31)
which are the parity of number of type-$1$ edge between $x_{m}$ and $x_{\ell}$, and the distance between the boundary vertices $k_{m}$ and $k_{\ell}$ along the facial walk $f$, respectively. To simplify the notation, let us put $f\cap\delta X=\{0,1,\dots,q-1\}$. We set the identity matrix, the weighted cyclic permutation matrix on $\mathbb{C}^{\{0,\dots,q-1\}}$ by
@@
From this expression, if $f\cap\delta X=\{0,\dots,q-1\}$, we have
$(S_{f})_{j,i}=\begin{cases}\displaystyle bc\;\frac{a^{j-i-1}\omega^{{\rm dist}_{f}(j,i)}}{1-a^{|f\cap\delta X|}}\times(-1)^{\bowtie_{f}(j,i)}&\text{: $i\neq j$,}\\ \\ \displaystyle bc\;\frac{a^{|f\cap\delta X|-1}\omega^{|f|}}{1-a^{|f\cap\delta X|}}+d&\text{: $i=j$.}\end{cases}$
@@
Pick up a facial walk $f=(\xi_{0},e_{0},\dots,\xi_{\kappa-1},e_{\kappa-1})\in\delta F$, and recall that $f\cap\delta X=\{x_{j_{0}},\dots,x_{j_{q-1}}\}$ and $f\cap\delta A_{is}=\{\xi_{k_{0}},\dots,\xi_{k_{q}}\}$. Here $1\leq q\leq\kappa$. Note that the $q$ tails interfere with the facial walk $f$. The tail incident to the boundary $x_{k_{i}}$ is denoted by $\mathrm{Tail}_{i}$. Let us give the inflow $1$ from the $\mathrm{Tail}_{i}$, and consider the outflow to the tail $\mathrm{Tail}_{j}$ ($i,j\in\{0,1,\dots,q-1\}$). Let us see that the $(j,i)$ element of the scattering matrix $S$ can be calculated according to the number of “nights” that a quantum walk stays at the face $f$ with the boundary vertices $x_{k_{i}}$ and $x_{k_{j}}$ from the time the quantum walk enters through entrance $i$ to the time the quantum walk leaves through exit $j$.
Consider the case for $i=0$. Let $t^{in}_{j}\in\delta A_{pr}^{+}$ and $t^{out}_{j}\in\delta A_{pr}^{-}$ be the arcs of $\mathrm{Tail}_{j}$ incident to the vertex $x_{k_{j}}\in\delta X$ (see Figure 11). Let “the day trip walk” from $\mathrm{Tail}_{0}$ to $\mathrm{Tail}_{j}$ be defined by
@@
$(t_{0}^{+},\boldsymbol{w}_{0},\overbrace{\xi_{k_{j}}^{+},e_{k_{j}},\dots,\xi_{k_{j+1}}^{-},\xi_{k_{j+1}}^{+},e_{k_{j+1}},\dots,\xi_{k_{j-1}}^{-},\xi_{k_{j-1}}^{+},e_{k_{j-1}},\dots,\xi_{k_{j}}^{-},}^{\text{$r$ times (nights)}}t_{j}^{-})$
@@
the weights associated with moving from the tail $t_{0}^{+}\in\delta A_{pr}^{+}$ to the quay $\xi_{k_{0}}^{+}\in A_{is}$, and from the quay $\xi_{k_{j}}^{-}\in A_{is}$ to the tail $t_{j}^{-}\in\delta A_{pr}^{-}$ are $c$ and $b$, respectively; the weight associated with moving from $\xi_{k_{s}}^{-}\in A_{is}$ to $\xi_{k_{s}}^{+}\in A_{is}$ is $a$. Remark that the weight of the closed path starting from $\xi_{{k_{0}}}^{+}$ and returning back to the same arc $\xi_{k_{0}}^{+}$ along the boundary face $f$ is $a^{q}\omega^{\kappa}$ since $(-1)^{\int_{f}\tau}=1$. Then set $\boldsymbol{s}^{(r)}(\cdot,0)\in\mathbb{C}^{\{0,1,\dots,q-1\}}$ ($r=0,1,\dots$) as the weight of $r$-night walk in the stationary state by $\boldsymbol{s}^{(r)}(j,0)=(a^{q}\omega^{\kappa})^{r}\times\boldsymbol{s}^{(0)}(j,0)$, $(r=1,\dots,\kappa-1)$ with
$\boldsymbol{s}^{(0)}(j,0)=\begin{cases}b\;c\;a^{j-1}\;\omega^{{\rm dist}_{f}(j,0)}\;(-1)^{\bowtie_{f}(j,0)}&\text{: $j\neq 0$}\\ b\;c\;a^{q-1}\;\omega^{\kappa}\;&\text{: $j=0$}\end{cases}$
@@
$(S)_{k_{j},k_{0}}$ $=\sum_{r=0}^{\infty}\boldsymbol{s}^{(r)}(j,0)=\begin{cases}bc\frac{a^{j-1}\omega^{k_{j}-k_{0}}}{1-a^{q}\omega^{\kappa}}\times(-1)^{\bowtie_{f}(j,0)}&\text{: $j\neq 0$}\\ \\ bc\frac{a^{q-1}\omega^{\kappa}}{1-a^{q}\omega^{\kappa}}+d&\text{: $j=0$}\end{cases}$
@@
$(S)_{k_{j},k_{i}}=\begin{cases}\displaystyle bc\;\frac{a^{j-i-1}\omega^{{\rm dist}_{f}(j,i)}}{1-a^{q}\omega^{\kappa}}\times(-1)^{\bowtie_{f}(j,i)}&\text{: $i\neq j$,}\\ \\ \displaystyle bc\;\frac{a^{q-1}\omega^{\kappa}}{1-a^{q}\omega^{\kappa}}+d&\text{: $i=j$,}\end{cases}$ (6.33)
@@
Let $(G,\rho_{1},\tau_{1})$ and $(G,\rho_{2},\tau_{2})$ be the rotation systems which are equivalent to each other. Let $S_{1}$ and $S_{2}$ be the scattering matrices of those resulting embbedings, respectively.
@@
Let $S_{1}$ and $S_{2}$ be the above. Recall that $\Delta(x)$ is defined in (5.24). Then we have
@@ 6.2 Detection of the orientability
Let us estimate whether the underlying closed surface is orientable or not by observing the outflow of the internal graph to an inflow under the Assumption 1.
By the hedgehog assignment of the tails in (6.34), $\delta X$ is isomorphic to $A_{br}$ by the following one-to-one correspondance $\varphi$ between $A_{br}$ and $\delta X$:
$\varphi(e)=t(\mathrm{is}^{\sharp}(\bar{e}))$ (6.36)
for any $e\in A_{br}$ (see Figure 11).
Here the definitions of $\xi^{-}$ and $\xi^{+}$ are defined in Definition 4.2. We label the scattering matrix by $e\in A_{br}$ by using the bijection map $\varphi$ from now on. Let us introduce the following subset of $A_{br}$. For any islands of $x\in X^{\tau}$, we define
$\delta A(x):=\{e\in A_{br}\;|\;t(e)\in t(A_{is}(x))\},$
which corresponds to the set of the vertices having a tail in the island $x$.
For example, let us insert $(+1)$-inflow into a single tail. The following theorem tells us that once we find an island emitting outflows whose signatures are not invariant, the underlying closed surface must be unorientable.
@@
For any islands of $x,y\in X^{\tau}$ with $x\neq y$, let the submatrix of $S$ formed from rows $\delta A(x)$ and columns $\delta A(y)$ be denoted by
$S|_{x,y}:=[(S)_{\varphi(e),\varphi(e^{\prime})}]_{e\in\delta A(x),\;e^{\prime}\in\delta A(y)}.$
Then under Assumption 1 with $a>0$, the underlying surface is orientable if and only if the signature of all the elements of $(S|_{x,y})_{\varphi(e),\varphi(e^{\prime})}$’s with $(S|_{x,y})_{\varphi(e),\varphi(e^{\prime})}\neq 0$ are invariant.
@@
$(S)_{\varphi(e),\varphi(e^{\prime})}/|S_{\varphi(e),\varphi(e^{\prime})}|=\begin{cases}(-1)^{\bowtie_{f}(t(e),t(e^{\prime}))}&\text{: $e$ and $e^{\prime}$ are included in the same face, }\\ 0&\text{: otherwise.}\end{cases}$
Here $\bowtie_{f}(z,z^{\prime})=\int_{z}^{z^{\prime}}\tau=\sum_{e\in f}\tau(e)=\sum_{e\in f\cap A_{br}}\tau(e)$ because $\tau(\xi)=0$ for any island arcs $\xi$. We remark that the resulting embedding of the rotation system $(G,\rho,\tau)$ is orientable if and only if $\int_{c}\tau=0$ for any closed paths in the rotation system $(G,\rho,\tau)$ by Operation in Section 3. This implies that if $\bowtie_{f}(z,z^{\prime})$ is independent of the choice of path, then $(G,\rho,\tau)$ is orientable. We also remark that for any rotation system $(G,\rho,\tau)$ whose resulting surface is orientable, there exists an equivalent rotation system $(G,\rho^{\prime},0)$ to it. Let us see that $\bowtie_{f}(z,z^{\prime})$ is independent of the choice of path if the underlying surface is orientable: assume that $(G,\rho,\tau)$ is orientable. If there exists two paths $p$ and $p^{\prime}$ with $o(p)=o(p^{\prime})$, $t(p)=t(p^{\prime})$ and $\int_{p}\tau\neq\int_{p}^{\prime}\tau$. However Lemma 5.1 implies that $\int_{p\cup\bar{p^{\prime}}}\tau=\int_{p}\tau-\int_{p^{\prime}}\tau=0$, which is contradiction. Therefore $\bowtie_{f}(t(e),t(e^{\prime}))$ is invariant under the choices from $e\in\delta A(x)$ and $e^{\prime}\in\delta A_{y}$ if and only if the underlying surface is orientable. ∎